OpenStudy (anonymous):

Find the domain of the function g(x)=4/6-7x

OpenStudy (anonymous):

looks like all real on my calc. it is just a line right?

OpenStudy (anonymous):

I know they are all real, but I do not know if it is x is not equal to 4, or x is greater than or equal to, or not equal to 6/7, or x is not equal to 0

OpenStudy (anonymous):

Those values of x for which the function results imaginary number will be rejected

OpenStudy (anonymous):

we can do like dis 4/6-7x=0 4-42x=0 x=4/42 hence the domain is R-{2/21} since this is the point where the domain becomes zero

OpenStudy (anonymous):

\[x \in R -\left\{ 6/7 \right\}\]